# Unsocial Intelligence: an Investigation of the Assumptions of AGI Discourse

### v3

@@ 3.2. Value-laden choices in AGI definitions: a taxonomy
3.2.1. Economic value
Some definitions of AGI are explicit about the values that they are centering. For example, OpenAI’s charter claims that “OpenAI’s mission is to ensure that artificial general intelligence (AGI)—by which we mean highly autonomous systems that outperform humans at most economically valuable work—benefits all of humanity” (OpenAI, 2018). This definition of AGI is meant to track the metric of performing economically valuable work. As others have pointed out (Morris et al., 2023), this assumes that other types of work are less valuable—a normative assumption.
@@ 3.2.2. Embodiment
In contrast, criteria like Wozniak’s coffee test require embodied AI systems. This test, first proposed by Apple founder Steve Wozniak (Fast Company, 2010), tasks the machine with going into an “average” American home and making coffee in it. It has since been included in definitions of AGI (Goertzel et al., 2012; Marcus, 2022). AI critics like Weizenbaum (1976) have also cited the field’s relative inattention to embodiment as a major shortcoming.
@@
3.2.3. Human-like processes versus outcomes
Another divisive choice in conceptualizing AGI or human-level AI is whether to insist on mechanisms or processes that mimic human cognitive processes. For example, Morris et al. (2023) want to prioritize outcomes (“what an AGI can accomplish”) over human-like processes because outcomes are more operationalizable. Excluding “processes” from the definition means that criteria based on the following are excluded: consciousness (Butlin et al., 2023; Lenharo, 2024; Searle, 1980; Summerfield, 2023; Smart, 2015), sentience (Schwitzgebel and Garza, 2015), and abilities modeled on human children (Turing, 2012; Nilsson, 2005; Gopnik, 2019, 2015; Summerfield, 2023).
@@ 3.2.3. Human-like processes versus outcomes
Rich Sutton’s “The Bitter Lesson” articulates this practical concern about maximizing return on investment (Sutton, 2019). Sutton frames this as a choice between focusing on implementing human-like processes versus focusing on “general methods that leverage computation”. He frames the choices as practically exclusive in the sense that “[t]ime spent on one is time not spent on the other” Sutton (2019). The bitter lesson is that building human-like processes into AI may be psychologically satisfying to researchers, but tends to be much less effective than focusing on “general purpose method[s]” that scale with computation.
@@
In short, different practitioners make different choices about what they consider to be “practical” for achieving AGI. Choices like these about effectiveness and how to best allocate limited resources are inherently value-laden—raising questions like for what goals, given whose beliefs and preferences, given what conditions? For example, Summerfield (2023) does not articulate what he means by “useful” or “effective”, let alone for whom it is useful or effective. Similarly, when Morris et al. (2023) argue for having shared, operationalizable (and thus outcome-based) standards for AGI on the grounds that the standards will be “useful”, they do not articulate whom it will be useful for, or consider whether it may be less than useful to some. Critics of current trends in AI have pointed to how its harms and benefits are distributed unevenly in ways that are correlated with existing power structures (Green, 2021; Abebe et al., 2020; Blodgett et al., 2020; Birhane and Guest, 2021; Costanza-Chock, 2020). Given this context about how AI is currently used, implicit assumptions that “useful” means “useful for everyone” need more justification.
3.2.4. Generality
@@
A point of widespread agreement among accounts that specifically favor the term “AGI” over “human-level intelligence” is the importance of generality (Goertzel, 2014; Morris et al., 2023; Agüera y Arcas and Norvig, 2023; Summerfield, 2023). Summerfield (2023) summarizes this as the view that a “generally intelligent individual is polymathic—good at everything.” [^4] Summerfield (2023) emphasizes the importance of the kind of flexibility that humans display in ordinary everyday activities—“whether performing daily rituals of basic hygiene, navigating the streets of their neighbourhood, negotiating the local market, deftly parenting their children, or judiciously managing household finances”—rather than in the kinds of abilities associated with “chess grandmasters” or “Nobel laureates”. This sense of generality is quite different from what gets emphasized by g. Rather than centering a putative variable that correlates with improved performance across all cognitive tasks, Summerfield emphasizes the multiple complex facets of the kind of cognitive flexibility that humans display. The recent popularization of the term “AGI” is often traced to the mid-2000s interest in contrasting “narrow AI” capable of “specific ‘intelligent’ behaviors in specific contexts” with systems that would be able to “self-adapt to changes in their goals or circumstances”, “generalize knowledge from one goal or context to others”, and so on (Goertzel, 2014). [^5] Goertzel (2014) traces the genealogy of the term “AGI” as follows. “The brief history of the term “Artificial General Intelligence” is as follows. In 2002, Cassio Pennachin and I were editing a book on approaches to powerful AI, with broad capabilities at the human level and beyond, and we were struggling for a title. I emailed a number of colleagues asking for suggestions. My former colleague Shane Legg came up with “Artificial General Intelligence,” which Cassio and I liked, and adopted for the title of our edited book [Goertzel et al. (2007)]. The term began to spread further when it was used in the context of the AGI conference series. A few years later, someone brought to my attention that a researcher named Mark Gubrud had used the term in a 1997 article on the future of technology and associated risks [Gubrud (1997)].”
@@
Although widespread, the strong emphasis on “generality, adaptability and flexibility” (Goertzel, 2014) in machine intelligence is not universal among accounts of human-level machine intelligence. For instance, Bostrom (2014) departs from this norm by denying that achieving human-level intelligence or superintelligence requires “general” intelligence at all. Bostrom sees this as part of his strong rejection of the requirement for human-like cognitive processes. He instead remains agnostic about the structure of abilities needed to match or outclass the instrumental performance of human intelligence.
@@
We argued earlier (Section 3.2.3) that assumptions about whether AGI should include human-like processes or not is value-laden because it is based on notions of “usefulness” or “effectiveness” that presuppose whom and what the assumptions are useful for. The same line of thought applies to arguments about whether the generality of abilities is important. Assumptions about what practical outcomes are desirable are needed to make the case for or against the importance of generality. Bostrom’s line of argument presumes that replicating the instrumental performance of human intelligence is the goal—but there are other possible goals for AI. Similarly, others’ assumption that being polymathic is essential to AGI glosses over two points: whether this is a desirable goal for AI, and the real diversity of human abilities (not all humans are similarly polymathic). Here we see an analogy with critiques of g on the basis that it ignores human diversity (see Section 2.1.1).
3.2.5. Individualism
@@
Conceptions of AGI also differ on whether they conceive of intelligence as a property of individuals—such as isolated humans or systems. One aspect of individualism about AGI is ontological, having to do with whether the entity ascribed “intelligence” is an individual. The psychometrics project of measuring and comparing the intelligence and cognitive skills of humans takes individuals as a key unit of analysis. AI evaluation practices like benchmarking, as currently practiced, mostly treat individual models as bearers of the properties they measure, and definers of AGI often propose tests that are mostly or entirely tests on individual agents (Chollet, 2019; Morris et al., 2023).
@@
By contrast, accounts like Bostrom (2014) and Attard-Frost (2023) reject ontological individualism. Bostrom argues that when thinking about what it means for machines to match or outclass human intelligence, the relevant unit of comparison includes not just the intelligence of an individual human but also “the combined intellectual capability of all of humanity” at present (Bostrom, 2014). Specifically, he argues that collectives—such as “firms, work teas, gossip networks, advocacy groups, academic communities, countries, even humankind as a whole” (Bostrom, 2014)—can be understood as a mechanism for increasing intelligence. Bostrom also refers to collective intelligence as “collective intellectual problem-solving capacity”. This rejection of ontological individualism is related to the longer tradition of thinking about collective intelligence (Lévy, 1994, 2010; Engelbart, 1962; Baltzersen, 2021; Suran et al., 2021; Araya and Marber, 2023; Landemore and Elster, 2012; Anderson, 2006; Putnam, 2011; Dewey, 2011; Peters and Heraud, 2015). Baltzersen (2021) proposes thinking of collective intelligence as “collective problem solving”, both in large and small groups of people. [^6] Lévy (2010) similarly defines collective intelligence as “the capacity of human collectives to engage in intellectual cooperation in order to create, innovate and invent”. In Section 5, we come back to a strand of this tradition that frames democracy as a “precondition for the full application of intelligence to the solution of social problems” (Putnam, 2011).
@@
A second dimension of individualism in conceptions of AGI is methodological: concerning whether measurements of intelligence are carried out in environments where individuals are acting by themselves or where agents interact in a social environment with other agents. Most AI benchmarks are focused on individual agents carrying out tasks by themselves. But an alternative vision is possible (Attard-Frost, 2023)—one that we describe in more detail in Section 4.1. These two different ways of measuring intelligence would lead to different prioritizations of resources for research and development, with attendant ethical implications. For example, if we measure intelligent agents’ performance in more social and interactive environments, would this also lead to AI systems having fewer unanticipated harmful effects when deployed into the “real world”? Arguably, many of the recent cases of unanticipated harmful effects from AI systems being deployed are due to these systems being evaluated pre-deployment on decontextualized tasks that ignore aspects of the social environment (Wolf et al., 2017; Ganguli et al., 2022; Saisubramanian et al., 2022).
3.2.6. Instrumentality Thesis
@@
A widespread assumption in discussions of AGI is what Russell (2019) calls the “standard model of intelligence”: “[Humans or] machines are intelligent to the extent that their actions can be expected to achieve their objectives.” Similarly, Legg and Hutter (2007) characterize intelligence as “measur[ing] an agent’s ability to achieve goals in a wide range of environments.” These accounts endorse what we call the instrumentality thesis: the view that intelligence is a means to whatever ends, final goals, or preferences an agent might have. [^7] Bostrom (2014) calls this the orthogonality thesis, which he defines as follows (Bostrom, 2014): “Intelligence and final goals are orthogonal: more or less any level of intelligence could in principle be combined with more or less any final goal.” We find it more helpful to highlight the similarity of this conception of intelligence with the conceptions of instrumental rationality that come from traditions like utility theory, or embrace Hume’s adage that “Reason is, and ought only to be the slave of the passions” (Hume, 1991; Schmitter, 2021). These accounts conceive of being more or less intelligent as categorically different from being better or worse at determining what matters, what goals are worth pursuing, or what is good for its own sake. Bostrom (2003) illustrates this feature of intelligence with the example of a paperclip maximizer: an agent with the final goal of building a world with maximally many paperclips. Paperclips are sometimes useful tools, such as for organizing paper documents. But paperclip maximizing as a final goal is patently not worthwhile. The instrumentality thesis is part of a common tradition in AI: thinking of intelligent agents as maximizing the expected utility of actions for satisfying given objectives or preferences.
@@ 3.2.7. Operationalizability/Measurability
Operationalizability may appear value-neutral at first glance. However, scholars have argued that simply having a common language and shared standards is in itself value-laden. Commensuration (the act of establishing common standards to measure things) allows large institutions to coordinate actions and automate decision-making, but less powerful groups often argue that some things cannot be measured in order to resist actions that can be justified through “rational choices” (Espeland and Stevens, 1998). Social scientists have also recognized that commensuration has performative effects, creating a kind of path dependency: “as commensuration gets built into practical organizations of labor and resources, it becomes more taken for granted and more constitutive of what it measures” (Espeland and Stevens, 1998). This type of path dependency has been recognized in various domains of commensuration, such as standardized grades of grain quality (Porter, 1996) and university rankings (Fowles et al., 2016). Previous work on ML benchmarking has also highlighted how benchmarking creates performativity and path dependency (Dehghani et al., 2021; Blili-Hamelin and Hancox-Li, 2023).
3.2.8. Choice of tasks/benchmarks
@@
Two prominent AI researchers recently argued that “[t]he most important parts of AGI have already been achieved by the current generation of advanced AI large language models” (Agüera y Arcas and Norvig, 2023). While these researchers define AGI roughly as instructable systems able to operate over a wide variety of topics, tasks, modalities and languages, the fact that the strengths of current LLMs are considered to already cover “the most important parts of AGI” implies that the things that LLMs cannot do are the less important parts of AGI. These LLM weaknesses include (as they acknowledge themselves) arithmetic and adhering to facts. However, they do not explain why these areas of reasoning are devalued relative to LLM strengths like translation or synthesizing code. Other researchers, in contrast, believe that arithmetical and logical reasoning are necessary components of AGI (Marcus, 2023). Agüera y Arcas and Norvig (2023)’s assertion that the most important aspects of AGI are achieved by current-day LLMs is value-laden. Choices of tasks and benchmarks are a key site of value-laden design choices (Blili-Hamelin and Hancox-Li, 2023; Bommasani, 2022). Deciding to focus on LLMs over deductive reasoning systems has ethical implications because these systems will change our societies in different ways.  [^8] For another critique of the reliance on benchmarks in relation to AGI, see Summerfield (2023).
3.2.9. Importing g into AI
@@
Firstly, the critiques of factor analysis as a methodology for finding a unique causal structure, as applied to the case of g in human intelligence (see Section 2.1.2), would also transfer over to the case of AI. Secondly, taking AI to have a directly analogous g factor means that all the value-laden assumptions around what human abilities count as important are imported over into the AGI case (see Section 2). For example, Chollet (2019), who also draws inspiration from g, takes for granted that there is a “space of tasks and domains that fit within the human experience” that we should use to measure intelligence—ignoring critiques of similar assumptions in the human intelligence literature (see Section 2.1.1).
3.3. Are deflationary views truly deflationary?

### v4

@@ 3.2 Value-laden choices in AGI definitions: a taxonomy
Economic value
Some definitions of AGI are explicit about the values that they are centering. For example, OpenAI’s charter claims that “OpenAI’s mission is to ensure that artificial general intelligence (AGI)—by which we mean highly autonomous systems that outperform humans at most economically valuable work—benefits all of humanity” (OpenAI 2018). This definition of AGI is meant to track the metric of performing economically valuable work. As others have pointed out (Morris et al. 2023), this assumes that other types of work are less valuable—a normative assumption.
@@
In contrast, criteria like Wozniak’s coffee test require embodied AI systems. This test, first proposed by Apple founder Steve Wozniak (Fast Company 2010), tasks the machine with going into an “average” American home and making coffee in it. It has since been included in definitions of AGI (Goertzel, Iklé, and Wigmore 2012; Marcus 2022). AI critics like Weizenbaum (1976) have also cited the field’s relative inattention to embodiment as a major shortcoming.
@@
Human-like processes versus outcomes
Another divisive choice in conceptualizing AGI or human-level AI is whether to insist on mechanisms or processes that mimic human cognitive processes. For example, Morris et al. (2023) want to prioritize outcomes (“what an AGI can accomplish”) over human-like processes because outcomes are more operationalizable. Excluding “processes” from the definition means that criteria based on the following are excluded: consciousness (Butlin et al. 2023; Lenharo 2024; Searle 1980; Summerfield 2023; Smart 2015), sentience (Schwitzgebel and Garza 2015), and abilities modeled on human children (Turing 2012; Nilsson 2005; Gopnik 2019, 2015; Summerfield 2023).
@@
Rich Sutton’s “The Bitter Lesson” articulates this practical concern about maximizing return on investment (Sutton 2019). Sutton frames this as a choice between focusing on implementing human-like processes versus focusing on “general methods that leverage computation”. He frames the choices as practically exclusive in the sense that “[t]ime spent on one is time not spent on the other” Sutton (2019). The bitter lesson is that building human-like processes into AI may be psychologically satisfying to researchers, but tends to be much less effective than focusing on “general purpose method[s]” that scale with computation.
@@
In short, different practitioners make different choices about what they consider to be “practical” for achieving AGI. Choices like these about effectiveness and how to best allocate limited resources are inherently value-laden—raising questions like for what goals, given whose beliefs and preferences, given what conditions? For example, Summerfield (2023) does not articulate what he means by “useful” or “effective”, let alone for whom it is useful or effective. Similarly, when Morris et al. (2023) argue for having shared, operationalizable (and thus outcome-based) standards for AGI on the grounds that the standards will be “useful”, they do not articulate whom it will be useful for, or consider whether it may be less than useful to some. Critics of current trends in AI have pointed to how its harms and benefits are distributed unevenly in ways that are correlated with existing power structures (Green 2021; Abebe et al. 2020; Blodgett et al. 2020; Birhane and Guest 2021; Costanza-Chock 2020). Given this context about how AI is currently used, implicit assumptions that “useful” means “useful for everyone” need more justification.
Generality
@@
A point of widespread agreement among accounts that specifically favor the term “AGI” over “human-level intelligence” is the importance of generality (Goertzel 2014; Morris et al. 2023; Agüera y Arcas and Norvig 2023; Summerfield 2023). Summerfield (2023) summarizes this as the view that a “generally intelligent individual is polymathic—good at everything.” [^4] Summerfield (2023) emphasizes the importance of the kind of flexibility that humans display in ordinary everyday activities—“whether performing daily rituals of basic hygiene, navigating the streets of their neighbourhood, negotiating the local market, deftly parenting their children, or judiciously managing household finances”—rather than in the kinds of abilities associated with “chess grandmasters” or “Nobel laureates”. This sense of generality is quite different from what gets emphasized by g. Rather than centering a putative variable that correlates with improved performance across all cognitive tasks, Summerfield emphasizes the multiple complex facets of the kind of cognitive flexibility that humans display. The recent popularization of the term “AGI” is often traced to the mid-2000s interest in contrasting “narrow AI” capable of “specific ‘intelligent’ behaviors in specific contexts” with systems that would be able to “self-adapt to changes in their goals or circumstances”, “generalize knowledge from one goal or context to others”, and so on (Goertzel 2014). [^5] Goertzel (2014) traces the genealogy of the term “AGI” as follows. “The brief history of the term “Artificial General Intelligence” is as follows. In 2002, Cassio Pennachin and I were editing a book on approaches to powerful AI, with broad capabilities at the human level and beyond, and we were struggling for a title. I emailed a number of colleagues asking for suggestions. My former colleague Shane Legg came up with “Artificial General Intelligence,” which Cassio and I liked, and adopted for the title of our edited book [Goertzel et al. (2007)]. The term began to spread further when it was used in the context of the AGI conference series. A few years later, someone brought to my attention that a researcher named Mark Gubrud had used the term in a 1997 article on the future of technology and associated risks [Gubrud (1997)].”
@@
Although widespread, the strong emphasis on “generality, adaptability and flexibility” (Goertzel 2014) in machine intelligence is not universal among accounts of human-level machine intelligence. For instance, Bostrom (2014) departs from this norm by denying that achieving human-level intelligence or superintelligence requires “general” intelligence at all. Bostrom sees this as part of his strong rejection of the requirement for human-like cognitive processes. He instead remains agnostic about the structure of abilities needed to match or outclass the instrumental performance of human intelligence.
@@
We argued earlier (Section 3.2) that assumptions about whether AGI should include human-like processes or not is value-laden because it is based on notions of “usefulness” or “effectiveness” that presuppose whom and what the assumptions are useful for. The same line of thought applies to arguments about whether the generality of abilities is important. Assumptions about what practical outcomes are desirable are needed to make the case for or against the importance of generality. Bostrom’s line of argument presumes that replicating the instrumental performance of human intelligence is the goal—but there are other possible goals for AI. Similarly, others’ assumption that being polymathic is essential to AGI glosses over two points: whether this is a desirable goal for AI, and the real diversity of human abilities (not all humans are similarly polymathic). Here we see an analogy with critiques of g on the basis that it ignores human diversity (see Section 2.1).
Individualism
@@
Conceptions of AGI also differ on whether they conceive of intelligence as a property of individuals—such as isolated humans or systems. One aspect of individualism about AGI is ontological, having to do with whether the entity ascribed “intelligence” is an individual. The psychometrics project of measuring and comparing the intelligence and cognitive skills of humans takes individuals as a key unit of analysis. AI evaluation practices like benchmarking, as currently practiced, mostly treat individual models as bearers of the properties they measure, and definers of AGI often propose tests that are mostly or entirely tests on individual agents (Chollet 2019; Morris et al. 2023).
@@
By contrast, accounts like Bostrom (2014) and Attard-Frost (2023) reject ontological individualism. Bostrom argues that when thinking about what it means for machines to match or outclass human intelligence, the relevant unit of comparison includes not just the intelligence of an individual human but also “the combined intellectual capability of all of humanity” at present (Bostrom 2014). Specifically, he argues that collectives—such as “firms, work teas, gossip networks, advocacy groups, academic communities, countries, even humankind as a whole” (Bostrom 2014)—can be understood as a mechanism for increasing intelligence. Bostrom also refers to collective intelligence as “collective intellectual problem-solving capacity”. This rejection of ontological individualism is related to the longer tradition of thinking about collective intelligence (Lévy 1994, 2010; Engelbart 1962; Baltzersen 2021; Suran, Pattanaik, and Draheim 2021; Araya and Marber 2023; Landemore and Elster 2012; Anderson 2006; Putnam 2011; Dewey 2011; Peters and Heraud 2015). Baltzersen (2021) proposes thinking of collective intelligence as “collective problem solving”, both in large and small groups of people. [^6] Lévy (2010) similarly defines collective intelligence as “the capacity of human collectives to engage in intellectual cooperation in order to create, innovate and invent”. In Section 5, we come back to a strand of this tradition that frames democracy as a “precondition for the full application of intelligence to the solution of social problems” (Putnam 2011).
@@
A second dimension of individualism in conceptions of AGI is methodological: concerning whether measurements of intelligence are carried out in environments where individuals are acting by themselves or where agents interact in a social environment with other agents. Most AI benchmarks are focused on individual agents carrying out tasks by themselves. But an alternative vision is possible (Attard-Frost 2023)—one that we describe in more detail in Section 4.1. These two different ways of measuring intelligence would lead to different prioritizations of resources for research and development, with attendant ethical implications. For example, if we measure intelligent agents’ performance in more social and interactive environments, would this also lead to AI systems having fewer unanticipated harmful effects when deployed into the “real world”? Arguably, many of the recent cases of unanticipated harmful effects from AI systems being deployed are due to these systems being evaluated pre-deployment on decontextualized tasks that ignore aspects of the social environment (Wolf, Miller, and Grodzinsky 2017; Ganguli et al. 2022; Saisubramanian, Zilberstein, and Kamar 2022).
Instrumentality Thesis
@@
A widespread assumption in discussions of AGI is what Russell (2019) calls the “standard model of intelligence”: “[Humans or] machines are intelligent to the extent that their actions can be expected to achieve their objectives.” Similarly, Legg and Hutter (2007) characterize intelligence as “measur[ing] an agent’s ability to achieve goals in a wide range of environments.” These accounts endorse what we call the instrumentality thesis: the view that intelligence is a means to whatever ends, final goals, or preferences an agent might have. [^7] Bostrom (2014) calls this the orthogonality thesis, which he defines as follows (Bostrom 2014): “Intelligence and final goals are orthogonal: more or less any level of intelligence could in principle be combined with more or less any final goal.” We find it more helpful to highlight the similarity of this conception of intelligence with the conceptions of instrumental rationality that come from traditions like utility theory, or embrace Hume’s adage that “Reason is, and ought only to be the slave of the passions” (Hume 1991; Schmitter 2021). These accounts conceive of being more or less intelligent as categorically different from being better or worse at determining what matters, what goals are worth pursuing, or what is good for its own sake. Bostrom (2003) illustrates this feature of intelligence with the example of a paperclip maximizer: an agent with the final goal of building a world with maximally many paperclips. Paperclips are sometimes useful tools, such as for organizing paper documents. But paperclip maximizing as a final goal is patently not worthwhile. The instrumentality thesis is part of a common tradition in AI: thinking of intelligent agents as maximizing the expected utility of actions for satisfying given objectives or preferences.
@@
Operationalizability may appear value-neutral at first glance. However, scholars have argued that simply having a common language and shared standards is in itself value-laden. Commensuration (the act of establishing common standards to measure things) allows large institutions to coordinate actions and automate decision-making, but less powerful groups often argue that some things cannot be measured in order to resist actions that can be justified through “rational choices” (Espeland and Stevens 1998). Social scientists have also recognized that commensuration has performative effects, creating a kind of path dependency: “as commensuration gets built into practical organizations of labor and resources, it becomes more taken for granted and more constitutive of what it measures” (Espeland and Stevens 1998). This type of path dependency has been recognized in various domains of commensuration, such as standardized grades of grain quality (Porter 1996) and university rankings (Fowles, Frederickson, and Koppell 2016). Previous work on ML benchmarking has also highlighted how benchmarking creates performativity and path dependency (Dehghani et al. 2021; Blili-Hamelin and Hancox-Li 2023).
Choice of tasks/benchmarks
@@
Two prominent AI researchers recently argued that “[t]he most important parts of AGI have already been achieved by the current generation of advanced AI large language models” (Agüera y Arcas and Norvig 2023). While these researchers define AGI roughly as instructable systems able to operate over a wide variety of topics, tasks, modalities and languages, the fact that the strengths of current LLMs are considered to already cover “the most important parts of AGI” implies that the things that LLMs cannot do are the less important parts of AGI. These LLM weaknesses include (as they acknowledge themselves) arithmetic and adhering to facts. However, they do not explain why these areas of reasoning are devalued relative to LLM strengths like translation or synthesizing code. Other researchers, in contrast, believe that arithmetical and logical reasoning are necessary components of AGI (Marcus 2023). Agüera y Arcas and Norvig (2023)’s assertion that the most important aspects of AGI are achieved by current-day LLMs is value-laden. Choices of tasks and benchmarks are a key site of value-laden design choices (Blili-Hamelin and Hancox-Li 2023; Bommasani 2022). Deciding to focus on LLMs over deductive reasoning systems has ethical implications because these systems will change our societies in different ways.  [^8] For another critique of the reliance on benchmarks in relation to AGI, see Summerfield (2023).
Importing g into AI
@@
Firstly, the critiques of factor analysis as a methodology for finding a unique causal structure, as applied to the case of g in human intelligence (see Section 2.1), would also transfer over to the case of AI. Secondly, taking AI to have a directly analogous g factor means that all the value-laden assumptions around what human abilities count as important are imported over into the AGI case (see Section 2). For example, Chollet (2019), who also draws inspiration from g, takes for granted that there is a “space of tasks and domains that fit within the human experience” that we should use to measure intelligence—ignoring critiques of similar assumptions in the human intelligence literature (see Section 2.1).
3.3 Are deflationary views truly deflationary?